OpenStudy (anonymous):

Given sin u = .315 and cos v = .759 and u and v are first quadrant angles. Find tan u and tan v using basic and pythagorean identities.

OpenStudy (anonymous):

First use this identity to get \(\cos\):\[ 1=\sin^2x+\cos^2x \implies \cos x=\sqrt{1-\sin^2 x} \]Then divide to get \(\tan\):\[ \tan(x) = \frac{\sin x}{\cos x} = \frac{\sin x}{\sqrt{1-\sin^2 x}} \]

OpenStudy (anonymous):

why do I try to find cos first?

OpenStudy (anonymous):

do I try to find cos and sin?

OpenStudy (anonymous):

and to find sin do I input sin v= sqrt 1-cos^2 ?